Circular No. 7313

Central Bureau for Astronomical Telegrams INTERNATIONAL ASTRONOMICAL UNION

Mailstop 18, Smithsonian Astrophysical Observatory, Cambridge, MA 02138, U.S.A.

IAUSUBS@CFA.HARVARD.EDU or FAX 617-495-7231 (subscriptions) BMARSDEN@CFA.HARVARD.EDU or DGREEN@CFA.HARVARD.EDU (science)

        URL http://cfa-www.harvard.edu/iau/cbat.html ISSN 0081-0304

Phone 617-495-7244/7440/7444 (for emergency use only)

2S 1417–624 M. H. Finger, Universities Space Research Association, reports for the

BATSE/CGRO team: “Hard x-ray pulses from the Be/x-ray transient pulsar 2S 1417–624 are currently being detected. The first detection was on Nov. 12. The average pulsed flux is 16 ± 3 mCrab (20–50 keV). After correction using the binary orbit of Finger et al. (1996, A.Ap. Supp. 120, 209), the pulse timing from Nov. 12–17 is consistent with a constant frequency of 56.99490(9) Hz. The last detection before the current outburst was on 1996 Dec. 21.”

COMET 10P/TEMPEL 2 C. W. Hergenrother and S. Larson, Lunar and Planetary Laboratory,

write: “Unfiltered CCD images of comet 10P taken with the Steward Observatory 2.3-m reflector by Hergenrother on Nov. 13.15 UT show evidence of shell-like jet-arc coma structure. Processing with a radial-gradiant filter shows multiple features spanning from p.a. 30^{o} clockwise through p.a. 100^{o} and extending out to 30′′ in the solar direction. Brightness enhancements are visible at p.a. 125^{o} and 345^{o}, the latter being near the predicted position angle of this comet’s traditional fan. Observations over a 1.5-hr span displayed no obvious changes. The coma morphology is similar to that reported in Nov. 1988 (Boehnhardt et al. 1990, Icarus 86, 58) and may be the result of viewing down the fan’s emission cone (Sekanina 1991, A.J. 102, 350). A processed image can be found at

http://www.lpl.arizona.edu/css/tempel2.html.”

LEONID METEORS 1999 A rather broad second peak in this year’s Leonid display appears to

have occurred ∼ Nov. 18.6–18.9 UT, in which visual counts of 100–300 meteors/hr were reported by observers in Hawaii (S. J. O’Meara), Japan (K. Mameda, Y. Kushida, R. Kushida, T. Kagawa; communicated via S. Nakano), and China (Rui Qi; communicated via J. Zhu).

V1493 AQUILAE CCD magnitude estimates, unfiltered unless otherwise noted: Oct.

15.743 UT, V = 15.73 ± 0.06 (M. Zejda, Brno, Czech Rep.); 31.730, V = 16.22 ± 0.04 (D. Hanžl, Brno, Czech Rep.); Nov. 1.726, 15.08 (G. Masi, Ceccano, Italy); 15.762, V = 16.1 ± 0.2 (Hanžl).

1999 November 20  \begin{matrix}c \\ ◯ \end{matrix} Copyright 1999 CBAT  Daniel W. E. Green